Trend Con is a silicon rubber designed to provide high clarity, superior mechanical properties, and short cycle time for high-speed manufacturers. One manufacturer used it to achieve smooth release from molds. The company’s projected growth would result in silicon costs of $26,000 next year and costs increasing by $2,000 per year through year 5. The interest rate is 10% per year.

a. What is the present worth of these costs?

i need to know the steps to enter all of this in my calculator. i dont understand how to do for example,(A/G, i%,N) i need the formula for all steps!!!

b. What is the future worth of this investment?
Answer #1

I = 10% = 0.1

PW of silicon costs = 26000*(P/A,10%,5) + 2000*(P/G,10%,5)

= 26000*[((1 + 0.1)^5-1)/(0.1*(1 + 0.1)^5)]+ 2000*[{((1 + 0.1)^5-1)/((0.1^2)(1 + 0.1)^5) - 5/(0.1*((1 + 0.1)^5))}]

= 26000*[((1.1)^5-1)/(0.1*(1.1)^5)]+ 2000*[{((1.1)^5-1)/((0.01)(1.1)^5) - 5/(0.1*((1.1)^5))}]

= 26000*[((1.61051-1)/(0.1*1.61051)]+ 2000*[{(1.61051-1)/((0.01)*1.61051) - 5/(0.1*(1.61051))}]

= 26000*[(0.61051)/(0.161051)]+ 2000*[{(0.61051)/(0.0161051) - 5/(0.161051)}]

= 26000*[3.7907867]+ 2000*[{37.9078676 - 31.04606615}]

= 26000*[3.7907867]+ 2000*[6.86180154]

= 98560.45600 + 13723.60308

= 112284.06

b. Future worth = 112284.06 * (1+0.1)^5 = 112284.06 * 1.61051 = 180834.60